OpenStudy (anonymous):

The principal P is borrowed at a simple interest rate r for a period of time t. Find the loan's future value A, or the total amount due at time t. Round answer to the nearest cent. P = $6,000, r = 6%, t = 2 years

OpenStudy (anonymous):

I=PRT I=6000*0.06*2 I=720 P+I=6720.00